STUDIES ON PLUTONIUM SPECIES

\underline{JEAN-PHILIPPE BLAUDEAU^{a}}, BRUCE E. BURSTEN, and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

We will report the status of our ongoing calculations on species containing plutonium. These have had a two-pronged approach: Density Functional Theory Methods (using ADF) studying the speciation of the various plutonium ions in water and multi-reference spin-orbit configuration interaction calculations (using COLUMBUS) studying the electronic structure of plutonyl ions. The speciation studies have concentrated on the comparison of coordination number and Pu-O bond lengths to experimental (from XANES and eXAFS) results. Recently, the electronic structure methods have allowed us to calculate transition spectra. The results for PuO_{2}^{2+} will be reported and compared to experiment.

aPresent Address: Ohio Supercomputer Center, ASC/MSRC 2435 5^{th} St. B676, Wright-Patterson AFB, Ohio, 45433